OpenStudy (anonymous):

find the real solutions of the equations √8x+48=x

OpenStudy (anonymous):

(√8x+48)=x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

8x+48=x^2?

OpenStudy (anonymous):

is the first x under the radical?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):